OpenStudy (anonymous):

two consecutive negativee integers have a product of 30. what are the integers? separate the answers with a comma.

OpenStudy (kainui):

So if the numbers are consecutive, then that means one is just +1 of the other. n is one of them. n+1 will be the other one. Now apparently their product is 30. So that just means multiply them together and you get 30 right? n*(n+1)=30 So now you can solve for n with this equation.

OpenStudy (anonymous):

ok

OpenStudy (kainui):

yep

OpenStudy (anonymous):

would it be -6,5

OpenStudy (kainui):

Almost, but -6*5=-30 which is wrong. See, when you solved for n that told you what possibilities you could have. n=-6 OR n=5. But remember these aren't both your numbers. The next number is n+1 right? So add 1 to both answers and you get: -6+1=-5 5+1=6 So now you see that: (-6)*(-5)=30 6*5=30 However in the question they said they wanted the consecutive negative integers, not the positive ones. So your answer is: -6 and -5.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the length and width of a rectangle are consecutive even integers. the area of the rectangle is 120 square units. what are the lenght and width of the rectangle

OpenStudy (anonymous):

would that be 10 ,12

OpenStudy (kainui):

Awesome, good job. =)